OpenStudy (anonymous):

**WILL MEDAL** Zuri made 2 spheres using plaster. The smaller sphere has a radius of 6 cm. The larger sphere has a radius of 24 cm. How much more plaster did Zuri use for the larger sphere? Use 3.14 to approximate pi and express your answer in hundredths. ___cm3 I got 56972.16

OpenStudy (anonymous):

@Michele_Laino ?

OpenStudy (michele_laino):

the volume V of a sphere whose radius is R, is given by the subsequent formula: \[\Large V = \frac{{4\pi }}{3}{R^3}\]

OpenStudy (michele_laino):

so we have: \[\Large V = \frac{{4\pi }}{3}{R^3} = \frac{{4 \cdot 3.14}}{3} \times {24^3} = ...?\]

OpenStudy (anonymous):

4.186*24^3 4.186*13824=57867.264

OpenStudy (anonymous):

Correct?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Wait so the answer is 57867.264?

OpenStudy (michele_laino):

I got this: 57876,48

OpenStudy (michele_laino):

I think that you got the right answer!

OpenStudy (anonymous):

Ok(:

OpenStudy (michele_laino):

thank you! :)